OpenStudy (anonymous):

What is the slope of the graph of 14x + 7y = –9?

OpenStudy (ajprincess):

try to solve the equation for y.

OpenStudy (ajprincess):

Follow these steps to solve. Subtract 14x both sides. Then divide both sides by 7. Does that help? @xXleeshaXx<3

OpenStudy (anonymous):

the form of all equations is y=mx+c,where ,m, is the slope and ,c, a aconstant now following the steps above gives y=(-14x-9)/7 y=-2x-9/7,now you can get the m=-2

OpenStudy (anonymous):